Sheffield, Inc. manufactures ergonomically designed computer furniture. Sheffield uses a job order costing system. On November 30, the Work in Process Inventory consisted of the following jobs ob No CC723 Computer caddy 25,000 CH291 Chair PS812 Printer stand Units Accumulated Cost $960,000 433,000 300,000 $1,693,000 18,000 26,000 On November 30, Sheffields Raw Materials Inventory account totaled $697,000, and its Finished Goods Inventory totaled $3,476,400. Sheffield applies manufacturing overhead on the basis of machine hours. The companys manufacturing overhead budget for the year totaled $5,000,000, and the company planned to use 1,000,000 machine hours during the year. Through the first eleven months of the year, the company used a total of 850,000 machine hours, total manufacturing overhead amounted to $4,284,500, and Cost of Goods Sold was $8,760,250. For the pursposes of this problem, ignore year-end disposition of over/under applied overhead Sheffield purchased $648,000 in raw materials in December and incurred the following costs for jobs in process that month: Materials Machine Direct Labor Direct Labor ob No. Issued CC723 CH291 PS812 $247,000 22,000 DS444 $260,000 17,000 Hours Hours Cost 13,200 4,000 14,300 16,900 $192,000 $18,500 13,000 4,600 $122,400 $47,200 $225,500 $138,000 The following jobs were completed in December and transferred to the Finished Goods Inventory: ob No. Item Units CC723 Computer caddy 25,000 CH291 Chair 18,000 DS444 Desk 9,000
(a) Calculate the total cost of each of the four jobs worked on in December CC723 CH291 PS812 DS444 1,924,400 728,700 1,872,500 1,248,000 Total Cost (b) Calculate the total manufacturing cost for December. Total manufacturing cost (c) Calculate the cost of goods manufactured for December Cost of goods manufactured (d) Calculate the balance in the Work in Process Inventory account on December 31 Balance in work in process e Assume that Sheffield sold 15,000 computer caddies, 12,000 chairs, and 4,500 desks in December. Calculate Cost of Goods Sold for the month of December and the ending Finished Goods Inventory balance on December 31. (For computation purpose round per unit rates to 2 decimal places, e.g. 3.54. Round final answers to O decmial places, e.g 45,000.) Cost of goods sold Finished goods inventory
Answer #1
a. Total cost of each job:
CC723 CH291 PS812 DS444
Accumulated cost 960000 433000 300000 1693000
Material cost 192000 18500 247000 260000
Direct labor cost 122400 47200 225500 138000
Overhead assinged (Note:1) 65000 23000 110000 85000
Total cost 1339400 521700 882500 2176000
Note:1
Predetermined overhead rate=Estimated manufacturing overhead/Estimated machune hours=5000000/1000000=$5 per machine hour
Overhead assigned=Machine hours worked*Predetermined overhead rate
CC723 CH291 PS812 DS444
Machine hours 13000 4600 22000 17000
Predetermined overhead rate 5 5 5 5
Overhead assigned 65000 23000 110000 85000
b. Total manufacturing cost=Material cost+direct labor cost+Overhead assigned
CC723 CH291 PS812 DS444 Total
Material cost 192000 18500 247000 260000 717500
Direct labor cost 122400 47200 225500 138000 533100
Overhead assinged 65000 23000 110000 85000 283000
Total manufacturing cost 1533600
c. Cost of goods manufactured:
$
Total manufacturing cost 1533600
Add: Beginning work in process inventory
(Accumulated cost at the beginning) 1693000
3226600
Less: Ending work in process inventory
(Refer d) 882500
Cost of goods manufactured 2344100
d. Balance in work in process inventory on Dec 31.
Balance in work in process consist of Jobs not completed.
Only uncompleted job is PS812
Hence,balance in work in process inventory on Dec 31=Cost of job PS812=$ 882500
e. CC723 CH291 DS444 Total
Total cost 1 1339400 521700 2176000
Units completed 2 25000 18000 9000
Cost per unit 3=1/2 53.58 28.98 241.78
Units sold 4 15000 12000 4500
Cost of goods sold 5=3*4 803640 347800 1088000 2239440
CC723 CH291 DS444 Total
Units completed 1 25000 18000 9000
Units sold 2 15000 12000 4500
Units in hand 3=1-2 10000 6000 4500
Cost per unit 4 53.58 28.98 241.78
Cost of finished goods inventory 5=3*4 535760 173900 1088000 1797660